# wavtokenizer: an efficient acoustic discrete codec tokenizer for audio language modeling

### v2

@@ Abstract
Language models have been effectively applied to modeling natural signals, such as images, video, speech, and audio. A crucial component of these models is the codec tokenizer, which compresses high-dimensional natural signals into lower-dimensional discrete tokens. In this paper, we introduce WavTokenizer, which offers several advantages over previous state-of-the-art (SOTA) acoustic codec models in the audio domain: 1) extreme compression. By compressing the layers of quantizers and the temporal dimension of the discrete codec, one-second audio of 24kHz sampling rate requires only a single quantizer with 40 or 75 tokens. 2) improved subjective reconstruction quality. Despite the reduced number of tokens, WavTokenizer achieves SOTA reconstruction quality with outstanding UTMOS scores and also inherently contains richer semantic information. Specifically, we achieve these results by designing a broader VQ space, extending contextual windows, improving attention networks, and introducing a powerful multi-scale discriminator and an inverse Fourier transform structure. We conduct extensive reconstruction experiments in the domains of speech, audio, and music. WavTokenizer exhibits competitive to superior performance across various objective and subjective metrics compared to SOTA models. We also evaluate WavTokenizer on semantic representation, VQ utilization, and adaptability to generative models. Comprehensive ablation studies confirm the necessity of each module in WavTokenizer. The code is available at https://github.com/jishengpeng/WavTokenizer.
@@ 1 Introduction
Recently, significant achievements have been made by large language models (LLMs) (Brown et al., 2020) in audio generative tasks, including multiple-speaker speech syntheses (Wang et al., 2023; Kharitonov et al., 2023; Jiang et al., 2023; 2024; Ji et al., 2024c), music generation (Agostinelli et al., 2023), and audio generation (Kreuk et al., 2022). Furthermore, the integration of the speech modality into unified large multimodal models also has garnered significant attention, such as SpeechGPT (Zhang et al., 2023a), AnyGPT (Zhan et al., 2024), GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, and Moshi (Défossez et al., 2024). These successes can be largely attributed to the utilization of discrete acoustic codec representations produced by neural codec models (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023). These discrete acoustic codec models bridge the gap between continuous speech signal and discrete-token-based language models, by discretizing high-rate audio signals into a finite set of tokens, hence enabling the application of LLM architectures to audio data.
@@ 4.1 Experimental Setup
Evaluation metrics and Implementation Details For objective evaluation of discrete codec models, following Vocos (Siuzdak, 2023), we employ the UTMOS (Saeki et al., 2022) automatic Mean Opinion Score (MOS) prediction system. UTMOS can yield scores highly correlated with human evaluations, closer to human perception than PESQ (Perceptual Evaluation of Speech Quality) (Rix et al., 2001), but it is restricted to 16 kHz sample rate. we also adopt the metrics in speech enhancement fields, such as PESQ, STOI (Short-time Objective Intelligibility), and the F1 score for voiced/unvoiced classification (V/UV F1). In addition to these objective metrics, following Encodec (Défossez et al., 2022), we also employ the subjective MUSHRA evaluation to assess the reconstruction performance of the codec. We employ the common subjective CMOS evaluation metrics to assess the performance of the downstream TTS model with the codec models. Details of the subjective evaluations are in Appendix C. Training and Inference Settings are detailed in Appendix A.
@@
Model Bandwidth $\downarrow$ Nq $\downarrow$ token/s $\downarrow$ UTMOS $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ V/UV F1 $\uparrow$ LibriTTS test-clean GT - - - 4.0562 - - - DAC 9.0kpbs 9 900 3.9097 3.9082 0.9699 0.9781 Encodec 6.0kbps 8 600 3.0399 2.7202 0.9391 0.9527 Vocos 6.0kbps 8 600 3.6954 2.8069 0.9426 0.9437 SpeechTokenizer 6.0kpbs 8 600 3.8794 2.6121 0.9165 0.9495 DAC 4.0kbps 4 400 3.4329 2.7378 0.9280 0.9572 HiFi-Codec 3.0kbps 4 400 3.7529 2.9611 0.9405 0.9617 HiFi-Codec 4.0kbps 4 300 3.9035 3.0116 0.9446 0.9576 Encodec 3.0kbps 4 300 2.3070 2.0517 0.9007 0.9198 Vocos 3.0kbps 4 300 3.5390 2.4026 0.9231 0.9358 SpeechTokenizer 3.0kbps 4 300 3.5632 1.9311 0.8778 0.9273 DAC 1.0kbps 1 100 1.4940 1.2464 0.7706 0.7941 WavTokenizer 0.5kbps 1 40 3.6016 1.7027 0.8615 0.9173 WavTokenizer 0.9kbps 1 75 4.0486 2.3730 0.9139 0.9382 LibriTTS test-other GT - - - 3.4831 - - - DAC 9.0kpbs 9 900 3.3566 3.7595 0.9576 0.9696 Encodec 6.0kbps 8 600 2.6568 2.6818 0.9241 0.9338 Vocos 6.0kbps 8 600 3.1956 2.5590 0.9209 0.9202 SpeechTokenizer 6.0kpbs 8 600 3.2851 2.3269 0.8811 0.9205 DAC 4.0kbps 4 400 2.9448 2.5948 0.9083 0.9404 HiFi-Codec 4.0kbps 4 400 3.0750 2.5536 0.9126 0.9387 HiFi-Codec 3.0kbps 4 300 3.3034 2.6083 0.9166 0.9318 Encodec 3.0kbps 4 300 2.0883 2.0520 0.8835 0.8926 Vocos 3.0kbps 4 300 3.0558 2.1933 0.8967 0.9051 SpeechTokenizer 3.0kbps 4 300 3.0183 1.7373 0.8371 0.8907 DAC 1.0kbps 1 100 1.4986 1.2454 0.7505 0.7775 WavTokenizer 0.5kbps 1 40 3.0545 1.6622 0.8336 0.8953 WavTokenizer 0.9kbps 1 75 3.4312 2.2614 0.8907 0.9172 LJSpeech GT - - - 4.3794 - - - DAC 9.0kpbs 9 900 4.3007 3.9022 0.9733 0.9757 Encodec 6.0kbps 8 600 3.2286 2.6633 0.9441 0.9555 Vocos 6.0kbps 8 600 4.0332 2.9258 0.9497 0.9459 SpeechTokenizer 6.0kpbs 8 600 4.2373 2.6413 0.9316 0.9452 DAC 4.0kbps 4 400 3.8109 2.7616 0.9338 0.9524 HiFi-Codec 4.0kbps 4 400 4.1656 2.7629 0.9446 0.9497 HiFi-Codec 3.0kbps 4 300 4.2692 2.9091 0.9485 0.9469 Encodec 3.0kbps 4 300 2.3905 2.0194 0.9058 0.9326 Vocos 3.0kbps 4 300 3.7880 2.5006 0.9310 0.9388 SpeechTokenizer 3.0kbps 4 300 3.9908 2.0458 0.9021 0.9299 DAC 1.0kbps 1 100 1.4438 1.2084 0.7822 0.8095 WavTokenizer 0.5kbps 1 40 4.0186 2.1142 0.9093 0.9406 WavTokenizer 0.9kbps 1 75 4.2580 2.4923 0.9312 0.9397
@@ 4.2 Main Results
Evaluation on Semantic Representation. We evaluate the semantic richness of different codec models on the ARCH benchmark (La Quatra et al., 2024). Notably, we opt not to utilize the conventional Superb benchmark (Yang et al., 2021) due to its exclusive focus on the speech domain, while ARCH enables further assessment of a Codec model in music and audio realms. The ARCH benchmark comprises 12 datasets in speech, music, audio domains (details in Appendix B). We extract embeddings corresponding to the discrete codebooks of an acoustic codec model as its respective representations and evaluate the classification accuracy of the codec model on ARCH datasets using its representations. For fair comparisons, we evaluate Encodec and DAC models on semantic representation, as they are under the same paradigm as WavTokenizer. The experimental results, as shown in Table 3, demonstrate that WavTokenizer substantially outperforms DAC and Encodec configured with a single quantizer or two quantizers on classification accuracy. Remarkably, on the AM and SLURP datasets in the speech domain, the MTT and IRMAS datasets in the music domain, and the FSD50K and VIVAE datasets in the audio domain, WavTokenizer surpasses DAC with nine quantizers and Encodec with eight quantizers on classification performance. These findings validate that WavTokenizer effectively captures rich semantic information.
@@
The results are shown in Table 4. In terms of audio quality (CMOS-Q) and audio prosody (CMOS-P), the speech synthesis model trained on WavTokenizer’s single-layer quantizer representations outperforms that trained on DAC’s 9-layer quantizer representations. This demonstrates that: 1) Acoustic models with a single-layer quantizer show significant potential in downstream autoregressive generative models, and 2) speech synthesis models using large-codebook acoustic representations can also synthesize high-quality audio, suggesting that large-codebook acoustic spaces have the potential to model speech as a special form of text. We will validate these findings on large multimodal models (Zhan et al., 2024) based on acoustic codec models that are trained on larger datasets in the future work.
Table 4: The subjective evaluations of various acoustic codec models for downstream speech synthesis models. GT denotes ground truth waveforms.
@@ Appendix F Limitation and Future work
While WavTokenizer is capable of reconstructing high-quality audio using only 75 tokens and demonstrates the potential of a single-layer quantizer with a large codebook space in downstream generative models, current acoustic codec models lack the understanding capabilities (ASR) found in semantic models (Hsu et al., 2021). This limitation constrains the development of codec models within unified multimodal understanding and generation frameworks, such as the GPT-4o paradigm. At present, distillation methods used in acoustic codec models serve more as temporary solutions. Unlike the powerful decoder we design in WavTokenizer, in future work, we aim to explore a more robust encoder module that can further improve compression, reconstruction, and semantic information retention.
@@
Appendix G acknowledgement
This work was supported by Alibaba Research Intern Program.

### v3

@@ Abstract
Language models have been effectively applied to modeling natural signals, such as images, video, speech, and audio. A crucial component of these models is the tokenizer, which compresses high-dimensional natural signals into lower-dimensional discrete tokens. In this paper, we introduce WavTokenizer, which offers several advantages over previous state-of-the-art (SOTA) acoustic codec models in the audio domain: 1) extreme compression. By compressing the layers of quantizers and the temporal dimension of the discrete codec, one-second audio of 24kHz sampling rate requires only a single quantizer with 40 or 75 tokens. 2) improved subjective reconstruction quality. Despite the reduced number of tokens, WavTokenizer achieves SOTA reconstruction quality with outstanding UTMOS scores and also inherently contains richer semantic information. Specifically, we achieve these results by designing a broader VQ space, extending contextual windows, improving attention networks, and introducing a powerful multi-scale discriminator and an inverse Fourier transform structure. We conduct extensive reconstruction experiments in the domains of speech, audio, and music. WavTokenizer exhibits competitive to superior performance across various objective and subjective metrics compared to SOTA models. We also evaluate WavTokenizer on semantic representation, VQ utilization, and adaptability to generative models. Comprehensive ablation studies confirm the necessity of each module in WavTokenizer.
@@ 1 Introduction
Recently, significant achievements have been made by large language models (LLMs) (Brown et al., 2020) in audio generative tasks, including multiple-speaker speech syntheses (Wang et al., 2023; Kharitonov et al., 2023; Jiang et al., 2023; 2024; Ji et al., 2024c), music generation (Agostinelli et al., 2023), and audio generation (Kreuk et al., 2022). Furthermore, the integration of the speech modality into unified large multimodal models also has garnered significant attention, such as SpeechGPT (Zhang et al., 2023a), AnyGPT (Zhan et al., 2024), GPT-4o, and Moshi (Défossez et al., 2024). These successes can be largely attributed to the utilization of discrete acoustic codec representations produced by neural codec models (Zeghidour et al., 2021; Défossez et al., 2022; Kumar et al., 2023). These discrete acoustic codec models bridge the gap between continuous speech signal and discrete-token-based language models, by discretizing high-rate audio signals into a finite set of tokens, hence enabling the application of LLM architectures to audio data.
@@ 4.1 Experimental Setup
Evaluation metrics and Implementation Details. For objective evaluation of discrete codec models, following Vocos (Siuzdak, 2023), we employ the UTMOS (Saeki et al., 2022) automatic Mean Opinion Score (MOS) prediction system. UTMOS can yield scores highly correlated with human evaluations, closer to human perception than PESQ (Perceptual Evaluation of Speech Quality) (Rix et al., 2001), but it is restricted to 16 kHz sample rate. we also adopt the metrics in speech enhancement fields, such as PESQ, STOI (Short-time Objective Intelligibility), and the F1 score for voiced/unvoiced classification (V/UV F1). In addition to these objective metrics, following Encodec (Défossez et al., 2022), we also employ the subjective MUSHRA evaluation to assess the reconstruction performance of the codec. We employ the common subjective CMOS evaluation metrics to assess the performance of the downstream TTS model with the codec models. Details of the subjective evaluations are in Appendix C. Training and Inference Settings are detailed in Appendix A.
@@
Dataset Model Bandwidth $\downarrow$ Nq $\downarrow$ token/s $\downarrow$ UTMOS $\uparrow$ PESQ $\uparrow$ STOI $\uparrow$ V/UV F1 $\uparrow$ LibriTTS test-clean GT - - - 4.0562 - - - DAC 9.0kpbs 9 900 3.9097 3.9082 0.9699 0.9781 Encodec 6.0kbps 8 600 3.0399 2.7202 0.9391 0.9527 Vocos 6.0kbps 8 600 3.6954 2.8069 0.9426 0.9437 SpeechTokenizer 6.0kpbs 8 600 3.8794 2.6121 0.9165 0.9495 DAC 4.0kbps 4 400 3.4329 2.7378 0.9280 0.9572 HiFi-Codec 3.0kbps 4 400 3.7529 2.9611 0.9405 0.9617 HiFi-Codec 4.0kbps 4 300 3.9035 3.0116 0.9446 0.9576 Encodec 3.0kbps 4 300 2.3070 2.0517 0.9007 0.9198 Vocos 3.0kbps 4 300 3.5390 2.4026 0.9231 0.9358 SpeechTokenizer 3.0kbps 4 300 3.5632 1.9311 0.8778 0.9273 DAC 1.0kbps 1 100 1.4940 1.2464 0.7706 0.7941 WavTokenizer 0.5kbps 1 40 3.6016 1.7027 0.8615 0.9173 WavTokenizer 0.9kbps 1 75 4.0486 2.3730 0.9139 0.9382 LibriTTS test-other GT - - - 3.4831 - - - DAC 9.0kpbs 9 900 3.3566 3.7595 0.9576 0.9696 Encodec 6.0kbps 8 600 2.6568 2.6818 0.9241 0.9338 Vocos 6.0kbps 8 600 3.1956 2.5590 0.9209 0.9202 SpeechTokenizer 6.0kpbs 8 600 3.2851 2.3269 0.8811 0.9205 DAC 4.0kbps 4 400 2.9448 2.5948 0.9083 0.9404 HiFi-Codec 4.0kbps 4 400 3.0750 2.5536 0.9126 0.9387 HiFi-Codec 3.0kbps 4 300 3.3034 2.6083 0.9166 0.9318 Encodec 3.0kbps 4 300 2.0883 2.0520 0.8835 0.8926 Vocos 3.0kbps 4 300 3.0558 2.1933 0.8967 0.9051 SpeechTokenizer 3.0kbps 4 300 3.0183 1.7373 0.8371 0.8907 DAC 1.0kbps 1 100 1.4986 1.2454 0.7505 0.7775 WavTokenizer 0.5kbps 1 40 3.0545 1.6622 0.8336 0.8953 WavTokenizer 0.9kbps 1 75 3.4312 2.2614 0.8907 0.9172 LJSpeech GT - - - 4.3794 - - - DAC 9.0kpbs 9 900 4.3007 3.9022 0.9733 0.9757 Encodec 6.0kbps 8 600 3.2286 2.6633 0.9441 0.9555 Vocos 6.0kbps 8 600 4.0332 2.9258 0.9497 0.9459 SpeechTokenizer 6.0kpbs 8 600 4.2373 2.6413 0.9316 0.9452 DAC 4.0kbps 4 400 3.8109 2.7616 0.9338 0.9524 HiFi-Codec 4.0kbps 4 400 4.1656 2.7629 0.9446 0.9497 HiFi-Codec 3.0kbps 4 300 4.2692 2.9091 0.9485 0.9469 Encodec 3.0kbps 4 300 2.3905 2.0194 0.9058 0.9326 Vocos 3.0kbps 4 300 3.7880 2.5006 0.9310 0.9388 SpeechTokenizer 3.0kbps 4 300 3.9908 2.0458 0.9021 0.9299 DAC 1.0kbps 1 100 1.4438 1.2084 0.7822 0.8095 WavTokenizer 0.5kbps 1 40 4.0186 2.1142 0.9093 0.9406 WavTokenizer 0.9kbps 1 75 4.2580 2.4923 0.9312 0.9397
@@ 4.2 Main Results
Evaluation on Semantic Representation. We evaluate the semantic richness of different codec models on the ARCH benchmark (La Quatra et al., 2024). Notably, we opt not to utilize the conventional Superb benchmark (Yang et al., 2021) due to its exclusive focus on the speech domain, while ARCH enables further assessment of a Codec model in music and audio realms. The ARCH benchmark comprises 12 datasets in speech, music, audio domains (details in Appendix B). We extract embeddings corresponding to the discrete codebooks of an acoustic codec model as its respective representations and evaluate the classification accuracy of the codec model on ARCH datasets using its representations. For fair comparisons, we evaluate Encodec and DAC models on semantic representation, as they are under the same paradigm as WavTokenizer. The experimental results, as shown in Table 3, demonstrate that WavTokenizer substantially outperforms DAC and Encodec configured with a single quantizer or two quantizers on classification accuracy. Remarkably, on the AM and SLURP datasets in the speech domain, the MTT and IRMAS datasets in the music domain, and the FSD50K and VIVAE datasets in the audio domain, WavTokenizer surpasses DAC with nine quantizers and Encodec with eight quantizers on classification performance.
@@
The results are shown in Table 4. In terms of audio quality (CMOS-Q) and audio prosody (CMOS-P), the speech synthesis model trained on WavTokenizer’s single-layer quantizer representations outperforms that trained on DAC’s 9-layer quantizer representations. This demonstrates that: 1) Acoustic models with a single-layer quantizer show significant potential in downstream autoregressive audio generation models, and 2) Speech synthesis models using large-codebook acoustic representations can also synthesize high-quality audio, suggesting that large-codebook acoustic spaces have the potential to model speech as a special form of text. We will validate these findings on large multimodal models (Zhan et al., 2024) based on acoustic codec models that are trained on larger datasets in future work.
Table 4: The Subjective Evaluations of various acoustic codec models for downstream text-to-speech synthesis models on the LibriTTS test set. GT denotes ground truth waveforms.
@@ 5 Conclusion
Acknowledgments
This work was done at Alibaba Group during the internship of Shengpeng Ji. I would like to express my sincere gratitude to my internship mentors, Qian Chen, Siqi Zheng, and Wen Wang, for their valuable guidance and support. This work was supported by the National Natural Science Foundation of China under Grant No.62222211.
@@ Appendix F Limitation and Future work
While WavTokenizer is capable of reconstructing high-quality audio using only 75 tokens and demonstrates the potential of a single-layer quantizer with a large codebook space in downstream generative models, current acoustic codec models lack the understanding capabilities (ASR) found in semantic models (Hsu et al., 2021). This limitation constrains the development of codec models within unified multimodal understanding and generation frameworks, such as the GPT-4o paradigm. At present, distillation methods used in acoustic codec models serve more as temporary solutions. Unlike the powerful decoder we design in WavTokenizer, in future work, we aim to explore a more robust encoder module that can further improve compression, reconstruction, and semantic information retention. Specifically, a key focus of our future work on WavTokenizer will be to explore elegant methods for significantly enhancing the semantic capacity of discrete tokens within the encoder.
@@
Appendix G WavTokenizer at 16kHz and 48kHz Sampling Rate
We have augmented the reconstruction evaluations by training both 16kHz and 48kHz versions of WavTokenizer using LibriTTS while maintaining a consistent downsampling factor of 320x. The hyperparameters are kept at their default settings. Experimental results on the LibriTTS test-clean dataset (4,837 samples) are presented in Table 10.
We find that WavTokenizer delivers high-quality and comparable audio reconstruction performance across different sampling rates of 16kHz, 24kHz, and 48kHz. Notably, at 16kHz sampling rate, WavTokenizer can effectively reconstruct audio with high fidelity using only 50 tokens. We also observed that all objective metrics, except for UTMOS, exhibit an increasing trend of increasing scores when the number of tokens increases. In contrast, UTMOS for WavTokenizer remains relatively stable. We hypothesize that the reconstruction quality of WavTokenizer should be stable across different sampling rates, to human perception of audio quality; hence, this observation of stable UTMOS aligns with prior finding Saeki et al. (2022) that UTMOS highly correlates with human subjective auditory perception.
Table 10: Objective reconstruction performance for WavTokenizer at different sampling rates on Librispeech test-clean. Sr denotes sampling rate. The results of WacTokenizer at 24kHz are compared to baseline models in Table 1.
Methods Sr Tokens/s UTMOS PESQ STOI V/UV F1
WavTokenizer 16000 50 3.9606 2.3240 0.9095 0.9375
WavTokenizer 24000 75 4.0486 2.3730 0.9139 0.9382
WavTokenizer 48000 150 3.9582 2.7230 0.9350 0.9496
Appendix H Reconstruction Speed
We evaluate the reconstruction speed of Semanticodec Encodec, DAC, and WavTokenizer on a single NVIDIA A100 80G GPU on the LibriTTS test-clean dataset. We calculate the real-time factor (RTF) by dividing the total reconstruction time by the duration of the generated audio. The results are shown in Table 11. Notably, despite incorporating the attention module in the decoder of WavTokenizer, its reconstruction speed remains remarkably fast. This can be attributed to two factors: (1) the use of a fast inverse Fourier transform, following Vocos, and (2) the low bitrate of WavTokenizer. These results demonstrate the high reconstruction efficiency of WavTokenizer.
Table 11: Reconstruction speed (measured by RTF) of different codec models on reconstruction on the LibriTTS test-clean dataset. RTF is computed by dividing the total reconstruction time by the duration of the generated audio.
Methods RTF
SemantiCodec 0.9483
Encodec 0.0175
DAC 0.0144
WavTokenizer 0.0098
Appendix I More TTS Results
Note that in Section 4.2, we present the subjective evaluation results of TTS models in the acoustic codec and LLM architecture in Table 4. We further supplement these results with synthesis accuracy and speaker similarity results. Building on the pre-trained TTS model using WavTokenizer as presented in Section 4.2, we further evaluate the synthesis accuracy, measured by WER, and speaker similarity SPK, computed by using WavLM to extract speaker embeddings for cosine similarity, on the zero-shot TTS task, following the same settings as the VALL-E-continue version (Wang et al., 2023). As shown in Table 12, in the audio codec and autoregressive LLM architecture for audio generation, with an identical generative model, the TTS model trained on WavTokenizer (0.9kbps) yields substantially lower WER than the TTS model trained on SOTA acoustic codec model DAC (9.0kbps), as 5.1% versus 6.9%, and also achieves better speaker similarity performance, as 0.61 versus DAC’s 0.59. This further confirms that WavTokenizer demonstrates excellent performance in the downstream TTS tasks.
Table 12: Word Error Rate (WER) and Speaker Similarity (SPK) of various acoustic codec models for downstream speech synthesis models. GT denotes ground truth waveforms. To evaluate Spk between the original prompt and the synthesized speech, we utilize the base-plus-sv version of WavLM.
[…]